OpenStudy (anonymous):

(y^2-16)/(y^2-25)*(y^2+2y-15)/(y^2-7y+12) i have to multiply help ASAP thanks

OpenStudy (anonymous):

\[\frac{\left(y^2-16\right) \left(y^2+2 y-15\right)}{\left(y^2-25\right) \left(y^2-7 y+12\right)}=\frac{(y-4) (y+4) (y+5) (y-3)}{(y-5) (y+5) (y-4) (y-3)}=\frac{y+4}{y-5} \]

OpenStudy (anonymous):

gracias

OpenStudy (anonymous):

@thania1 did you get it?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yups merci

OpenStudy (anonymous):

pas de quoi